OpenStudy (anonymous):

What number should be added to the expression x2 + 2x to change it into a perfect square trinomial? A. 1 B. 2 C. 3 D. 4

OpenStudy (anonymous):

Just complete the square if you want to find the last term that makes it a PST. ax^2 + bx + c = 0 x^2 + 4x Complete the square: take the "b" term and divide it by 2, then square that number and that will become your "c". x^2 + 4x + 4 (b = 4, 4/2 is 2 and 2^2 is 4). So your answer is 4.

OpenStudy (epoweritheta):

x^2+2x+1=(x+1)^2; so 1 should be added ; (x+a)^2=x^2+2ax+a^2 and compare coefficients

OpenStudy (anonymous):

Is @ikatouni correct?

OpenStudy (epoweritheta):

no i think he took the wrong question